OpenStudy (anonymous):

help solve using the elimination method 5u + 2u = -17 3u + v = -8

OpenStudy (callisto):

Is the question typed correctly? I mean for the first equation 5u + 2u = -17

OpenStudy (theviper):

multiply first by 3 & second by 5

OpenStudy (theviper):

it can be 5u+2v=-17

OpenStudy (anonymous):

the question should be 5u + 2v = -17

OpenStudy (theviper):

k then subtract first eq from 2:)

OpenStudy (theviper):

follow my or @Callisto step 1 by 1:)

OpenStudy (callisto):

Okay, the simpler way to do it is to multiply the second equation by 2. Then subtract (1) by (2) That is 5u + 2v = -17 -(1) 3u + v = -8 -(2) (2)x2 - (1) 2( 3u + v ) - (5u + 2v) = 2(-8) - (-17)

OpenStudy (theviper):

yes right

OpenStudy (theviper):

got it @marie52 ????

OpenStudy (anonymous):

yes! Thanks!

OpenStudy (theviper):

yw:)